In a survey, 15 people were asked how much they spent on their child's last birthday gift. The results were roughly bell-shaped with a mean of $40 and standard deviation of $7. Find the margin of error at a 95% confidence level.
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = $40

sample standard deviation = s = $7

sample size = n = 15

Degrees of freedom = df = n - 1 = 15-1= 14

At 95% confidence level the t is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,14 = 2.145

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.145 * (7 / \sqrt 15)

E = 3.9

The margin of error at a 95% confidence level is 3.9